OpenStudy (anonymous):

Solve M = 2x + 3y for y

OpenStudy (zrsmith):

M=2x+3y-2x (subtract 2x to get y by itself) -2x+M = 3y then divide both side by 3 -2/3*x+M/3 = y or y= -2/3*x+1/3*M